OpenStudy (anonymous):

6oz of mixed nuts containing 60% peanuts were mixed with 14oz of another kind of mixed nuts that contain 30% peanuts.peanuts are what % of the new mixture?

OpenStudy (amistre64):

6(.6) + 14(.3) = 20(P) 3.6 + 4.2 = 20(P) 7.8/20 = (P) P = .39 or 39%

OpenStudy (amistre64):

am I right?

OpenStudy (anonymous):

yes thank u!i get it

OpenStudy (amistre64):

yay!! :) youre welcome